1. What mass of CO2 is produced from burning 6.00 g C2H6 in excess oxygen? A. 11.7 g B. 8.80 g C. 17.6 g D. 4.40 g E. none of these

2. Consider the following unbalanced reaction: Al (s) + Br2 (l) --> Al2Br6 (s) How many moles of Br2 (l) are required to react with 10 moles of Al? A. 10 moles B. 60 moles C. 20 moles D. 30 moles E. 15 moles
Answer #1

1)

Molar mass of C2H6 = 2*MM(C) + 6*MM(H)

= 2*12.01 + 6*1.008

= 30.068 g/mol

mass of C2H6 = 6.0 g

we have below equation to be used:

number of mol of C2H6,

n = mass of C2H6/molar mass of C2H6

=(6.0 g)/(30.068 g/mol)

= 0.1995 mol

we have the Balanced chemical equation as:

2 C2H6 + 7 O2 ---> 4 CO2 + 6 H2O

Molar mass of CO2 = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

From balanced chemical reaction, we see that

when 2 mol of C2H6 reacts, 4 mol of CO2 is formed

mol of CO2 formed = (4/2)* moles of C2H6

= (4/2)*0.1995

= 0.3991 mol

we have below equation to be used:

mass of CO2 = number of mol * molar mass

= 0.3991*44.01

= 17.6 g

Answer: C

Only 1 question at a time please